# Dimension-independent convergence rates of randomized nets using median-of-means

### v1

@@ 4 Tractability under the completely random design
$\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\text{ for }j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
@@
$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
@@
Because each element of $\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
$\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
@@
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
$=$ $A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda},$
@@
$2^{-T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant B^{|u|}_{\alpha,\lambda}.$
@@
$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{-|v|}_{\alpha,\lambda}\Big{)}$ (27)
$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{-|v|}\Big{)}$
@@
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}.$ (29)
@@
$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}|K_{u}(T_{u,m})|\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{T_{u,m}/(\alpha+\lambda)}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
$=$ $\frac{\delta 2^{m}}{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
@@
Now we are ready to show the following error bounds on $\hat{\mu}^{(r)}_{\infty}$:
@@
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{2\theta(\alpha+\lambda)(m-4)}}$
@@
Our conclusion follows after plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and equation (30). ∎
@@
Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$. For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, $N=r2^{m}$ is the total number of function evaluations.
To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$. Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$ that
@@
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [22, Subsection 3.2].
@@
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=$ $1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})\Big{)},$
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ and a constant $C^{\prime}_{\eta}$ independent of $s$ [10, Lemma 3]. Substituting into equation (31), equation (33) holds with a constant $C_{\eta}$ independent of $s$, fulfilling the definition of strong tractability [20]. Notably, equation (34) resembles the tractability conditions assumed by [8, Corollary 3.13] for polynomial lattices to achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate.

### v2

@@ 4 Tractability under the completely random design
$\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\ \forall j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
@@
$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
@@
Because each element of $\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
$\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
@@
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
@@
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}.$
@@
$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda}\Big{)}$ (27)
$\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|}\Big{)}$
@@
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}.$ (29)
@@
$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}|K_{u}(T_{u,m})|\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{T_{u,m}/(\alpha+\lambda)}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
$=$ $\frac{\delta 2^{m}}{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
@@
Now we are ready to prove the following error bounds on $\hat{\mu}^{(r)}_{\infty}$:
@@
$2^{-m+4}\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{(2\theta(\alpha+\lambda)+1)m}}$
@@
Our conclusion follows by plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and equation (30). ∎
@@
Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$. For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, where $N=r2^{m}$ is the total number of function evaluations.
To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$. Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$,
@@
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [25, Subsection 3.2].
@@
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=$ $1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})\Big{)},$
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [11, Lemma 3]. Putting the above bounds into equation (31), we conclude that equation (33) holds with a constant $C_{\eta}$ independent of $s$. Notably, equation (34) resembles the tractability conditions assumed by [9, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.